# PCF Learned Sort: a Learning Augmented Sort Algorithm with $\mathcal{O}(n\log\log n)$ Expected Complexity

### v1

@@ 3.1 Method Overview
Let $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ be the array formed by sampling $n$ times independently according to $f_{\mathcal{I}}(x)$. Assume that there exist a model-based bucketing algorithm $\mathcal{M}$ and a constant $d~{}(\in(0,1))$ that satisfy the following for any interval $\mathcal{I}~{}(\subset\mathcal{D})$:
  • $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$, with an expected complexity of $\mathcal{O}(n)$. That is, $\mathcal{M}$ can take the array $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ as input and divide it into $\gamma+1$ buckets, $\{\bm{c}_{j}\}_{j=1}^{j=\gamma+1}$ satisfying Equation 5, with an expected complexity of $\mathcal{O}(n)$.
  • $\gamma+1=\mathcal{O}(n)$.
  • $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}\left(\frac{1}{\log n}\right)$.
The expected complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ is $\mathcal{O}(n\log\log n)$.
This lemma can be proved intuitively by the following two points: (i) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (ii) the expected total computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n)$. (i) is the same as in the explanation of the proof of Lemma 3.1. (ii) can be shown from the fact that the expected computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n\log n)$ with probability $\mathcal{O}(\frac{1}{\log n})$, and $\mathcal{O}(n)$ in other cases. See Section A.2 for the exact proof.
Note that the assumption of Lemma 3.3 includes “$\mathcal{M}$ works well with high probability for any $\mathcal{I}$.” This is because our Learned Sort algorithm recursively repeats the model-based bucketing. The range of elements in the bucket, i.e., the input array in the next recursion step, can be any interval $\mathcal{I}~{}(\subset\mathcal{D})$.
@@ 3.2 PCF Learned Sort
We propose PCF Learned Sort as an implementation that satisfies the assumptions of Lemma 3.1 and Lemma 3.3, and thus has a guarantee that the worst-case complexity is $\mathcal{O}(n\log n)$ and the expected complexity is $\mathcal{O}(n\log\log n)$. PCF Learned Sort approximates the CDF by a Piecewise Constant Function (PCF). Piecewise constant function is a function that has intervals of equal width and outputs the same value in each interval (the right side of Figure 1). The study that develops a Learned Index with a theoretical guarantee on its complexity (Zeighami & Shahabi, 2023) also used PCF as a CDF model.
The model-based bucketing method in PCF Learned Sort $\mathcal{M}_{\mathrm{PCF}}$ trains the CDF model $\tilde{F}$ in the following way. First, the parameters $\alpha\in\{1,\dots,n\}$ and $\beta\in\mathbb{N}$ are determined by $n$, the length of the input array. The parameter $\alpha$ is the number of samples to train the CDF model, and $\beta$ is the number of intervals in the PCF. Next, the piecewise constant function is trained by counting the number of samples in each interval. Using $x_{\mathrm{min}}=\min_{i}{x_{i}}$ and $x_{\mathrm{max}}=\max_{i}{x_{i}}$, let $i(x)$ be a function defined as follows:
$i(x)=\left\lfloor\frac{x-x_{\mathrm{min}}}{x_{\mathrm{max}}-x_{\mathrm{min}}}\beta\right\rfloor+1.$ (7)
$\alpha$ samples are taken at random from $\bm{x}$ to form $\bm{a}~{}(\in\mathcal{D}^{\alpha})$, and then $i(x)$ is used to form the array $\bm{b}~{}(\in{\mathbb{Z}_{\geq 0}}^{\beta+1})$ defined as follows:
$b_{i}=\left|\left\{j\in\{1,\dots,\alpha\}\mid i(a_{j})\leq i\right\}\right|.$ (8)
This counting is corresponding to the training of the PCF. Note that $\bm{b}$ is an non-decreasing non-negative array and $b_{\beta+1}=\alpha$, i.e., $0\leq b_{1}\leq b_{2}\leq\dots\leq b_{\beta+1}=\alpha$.
The inference for the CDF model $\tilde{F}(x)$ is performed using $i(x)$, $\bm{b}$, and the following equation:
$\tilde{F}(x)=\frac{b_{i(x)}}{\alpha}.$ (9)
Since $i(x)$ is a non-decreasing function and $\bm{b}$ is also a non-decreasing array, $\tilde{F}(x)$ is a non-decreasing function. Also, $0\leq\tilde{F}(x)\leq 1$ because $0\leq b_{i}\leq\alpha$ for every $i$.
@@
Let $\sigma_{1}$ and $\sigma_{2}$ be respectively the lower and upper bounds of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$. That is, $x\in\mathcal{D}\Rightarrow\sigma_{1}\leq f(x)\leq\sigma_{2}$.
Then, in model-based bucketing of $\bm{x}_{\mathcal{I}}~{}(\in\mathcal{I}^{n})$ to $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$ using $\mathcal{M}_{\mathrm{PCF}}$, the following holds for any interval $\mathcal{I}~{}(\subset\mathcal{D})$:
$\begin{split}&K\geq 1\\ \Rightarrow&\Pr[\exists j,|\bm{c}_{j}|>\delta]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\},\end{split}$ (10)
where
$K\coloneqq\frac{\gamma\delta}{2n}-\frac{2\sigma_{2}\gamma}{\sigma_{1}\beta}.$ (11)
The proof of this lemma is based on and combines proofs from two existing studies. The first is Lemma 5.2. from a study of IPS4(Axtmann et al., 2022), an efficient sample sort. This lemma guarantees the probability of a “successful recursion step” when selecting pivots from samples and using them to perform a partition. Although the definition of “successful recursion step” in this paper is similar to our definition of “successful bucketing” (more precisely, in the special case of $\delta=3n/\gamma$), this lemma is for the method that do not use the CDF model, so the proof cannot be applied directly to our case. Another proof we refer to is the proof of Lemma 4.5. from a study that addressed the computational complexity guarantee of the Learned Index (Zeighami & Shahabi, 2023). This lemma provides a probabilistic guarantee for the error between the output of the PCF and the empirical CDF. Some modifications are required to adapt it to the context of sorting and to attribute it to the probability of bucketing failure, i.e., $\Pr[\exists j,|\bm{c}_{j}|>\delta]$. By appropriately combining the proofs of these two lemmas, Lemma 3.4 can be proved. The exact proof is given in Section A.3.
Using Lemma 3.1, Lemma 3.3 and Lemma 3.4, we can prove the following theorems.
@@
If we use $\mathcal{M}_{\mathrm{PCF}}$ as the bucketing method and set $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$.
@@
When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$. This is because (i) the PCF is trained in $\mathcal{O}(\alpha+\beta)=\mathcal{O}(n^{3/4})$, and (ii) the total complexity of inference for $n$ elements is $\mathcal{O}(n)$, since the inference is performed in $\mathcal{O}(1)$ per element. Therefore, from Lemma 3.1, the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$. ∎
@@
Let $\sigma_{1}$ and $\sigma_{2}$ be the lower and upper bounds, respectively, of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$.
Then, if we use $\mathcal{M}_{\mathrm{PCF}}$ as the bucketing method and set $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$.
@@
When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$.
Since $K=\Omega(\sqrt{n})$ when $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, $K\geq 1$ for sufficiently large $n$, and
$\begin{split}\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}&=\mathcal{O}(n^{\frac{1}{4}}\exp(-\sqrt{n}))\\ &\leq\mathcal{O}\left(\frac{1}{\log n}\right).\end{split}$ (12)
Therefore, from Lemma 3.3, the expected computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$. ∎
Note that the exact value of $\sigma_{1}$ and $\sigma_{2}$ is not required to run PCF Learned Sort, since the parameters for this algorithm, i.e., $\alpha$, $\beta$, $\gamma$ and $\delta$, are determined without the knowledge. In other words, PCF Learned Sort can sort in expected $\mathcal{O}(n\log\log n)$ complexity as long as $0<\sigma_{1}\leq f(x)\leq\sigma_{2}<\infty$, even if it does not know the exact value of $\sigma_{1}$ and $\sigma_{2}$. If $\sigma_{1}$ and $\sigma_{2}$ do not satisfy the assumption of Theorem 3.6, i.e., $\sigma_{1}=0$ or $\sigma_{2}=\infty$, then the expected complexity of PCF Learned Sort is increased to $\mathcal{O}(n\log n)$, but from Theorem 3.5, it cannot be greater than $\mathcal{O}(n\log n)$.
@@ 4 Experiments
In this section, we confirm our theorems experimentally. First, we confirm that the computational complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ for both synthetic and real data. Then, we conduct experiments with various parameter settings and experimentally confirm Lemma 3.4, a lemma that bounds the probability of bucketing failure and is an important lemma to guarantee the expected computational complexity of PCF Learned Sort.
4.1 Results on Synthetic Datasets
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 2: Number of operations to sort the synthetic array. Below each graph is a histogram that visualizes the distribution of each dataset.

### v3

@@ 3.2 Learned-Sort Framework
Let $\bm{x}_{\mathcal{I}}~(\in\mathcal{I}^{n})$ be the array formed by sampling $n$ times independently according to $f_{\mathcal{I}}(x)$. Assume that there exist a model-based bucketing algorithm $\mathcal{M}$ and a constant $d~(\in(0,1))$ that satisfy the following three conditions for any interval $\mathcal{I}~(\subseteq\mathcal{D})$: (i) $\mathcal{M}$ can perform bucketing (including model training and inferences) on an array of length $n$, with an expected complexity of $\mathcal{O}(n)$, (ii) $\gamma+1=\mathcal{O}(n)$, and (iii) $\Pr[\exists j,|\bm{c}_{j}|\geq\lfloor n^{d}\rfloor]=\mathcal{O}\left(\frac{1}{\log n}\right)$. Also, assume that the standard sort has an expected complexity of $\mathcal{O}(n\log n)$. Then, the expected complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ is $\mathcal{O}(n\log\log n)$.
This lemma can be proved intuitively by the following two points: (1) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (2) the expected total computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n)$. The first point is the same as in the explanation of the proof of Lemma˜3.1. The second point holds because the expected computational complexity from the $i$-th to the $(i+1)$-th recursion depth is $\mathcal{O}(n\log n)$ with probability $\mathcal{O}(\frac{1}{\log n})$, and $\mathcal{O}(n)$ in other cases. See Section˜A.2 for the exact proof.
Note that the assumption of Lemma˜3.3 includes $\mathcal{M}$ works well with high probability for any $\mathcal{I}~(\subseteq\mathcal{D})$.” This is because our Learned Sort algorithm recursively repeats the model-based bucketing. The range of elements in the bucket, i.e., the input array in the next recursion step, can be any interval $\mathcal{I}~(\subseteq\mathcal{D})$.
@@ 3.3 PCF Learned Sort
We now instantiate the framework with PCF Learned Sort, which satisfies the assumptions of Lemma˜3.1 and Lemma˜3.3, thereby providing both worst-case and expected-time guarantees. PCF Learned Sort approximates the CDF using a Piecewise Constant Function (PCF) with $\beta$ equal-width bins; the model output is constant within each bin (see the right panel of Figure 1). The study that develops a Learned Index with a theoretical guarantee on its complexity (zeighami2023distribution) also used PCF as a CDF model. While our framework admits more expressive CDF models (such as spline-based or neural models), we focus on PCF due to its minimal training and inference cost and the tractability of its theoretical analysis.
The model-based bucketing method in PCF Learned Sort $\mathcal{M}_{\mathrm{PCF}}$ trains the CDF model $\tilde{F}$ as follows. The parameters $\alpha\in\{1,\dots,n\}$ and $\beta\in\mathbb{N}$ are determined by $n$, where $\alpha$ is the number of samples used to train the model and $\beta$ is the number of intervals in the PCF. The PCF is trained by counting the number of samples in each interval. We define $i(x)=\lfloor(x-x_{\min})\beta/(x_{\max}-x_{\min})\rfloor+1$, where $x_{\min}=\min_{i}x_{i}$ and $x_{\max}=\max_{i}x_{i}$. From $\bm{x}$, $\alpha$ samples are taken to form $\bm{a}\in\mathcal{D}^{\alpha}$, and $i(x)$ is used to construct $\bm{b}\in\mathbb{Z}_{\geq 0}^{\beta+1}$ with $b_{i}=\left|\left\{j\in\{1,\dots,\alpha\}\mid i(a_{j})\leq i\right\}\right|$. This counting procedure trains the PCF. Note that $\bm{b}$ is an non-decreasing non-negative array and $b_{\beta+1}=\alpha$, i.e., $0\leq b_{1}\leq b_{2}\leq\dots\leq b_{\beta+1}=\alpha$.
Inference for $\tilde{F}(x)$ is then given by $\tilde{F}(x)=\frac{b_{i(x)}}{\alpha}$. Since $i(x)$ and $\bm{b}$ are non-decreasing, $\tilde{F}(x)$ is also non-decreasing. Also, $0\leq\tilde{F}(x)\leq 1$ because $0\leq b_{i}\leq\alpha$ for every $i$.
@@
Let $\sigma_{1}$ and $\sigma_{2}$ be respectively the lower and upper bounds of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$. That is, $x\in\mathcal{D}\;\;\Rightarrow\;\;\sigma_{1}\leq f(x)\leq\sigma_{2}$. Then, in model-based bucketing of $\bm{x}_{\mathcal{I}}~(\in\mathcal{I}^{n})$ to $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$ using $\mathcal{M}_{\mathrm{PCF}}$, the following holds for any interval $\mathcal{I}~(\subseteq\mathcal{D})$:
$K\coloneqq\frac{\gamma\delta}{2n}-\frac{2\sigma_{2}\gamma}{\sigma_{1}\beta}\geq 1\;\;\Rightarrow\;\;\Pr[\exists j,|\bm{c}_{j}|>\delta]\leq\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}.$ (1)
The proof of this lemma is based on and combines proofs from two existing studies. The first is Lemma 5.2. from a study of IPS4(axtmann2022engineering), an efficient sample sort. This lemma guarantees the probability of a “successful recursion step” when selecting pivots from samples and using them to perform a partition. This lemma is for the method that does not use the CDF model, so the proof cannot be applied directly to our case. Another proof we refer to is the proof of Lemma 4.5. from a study that addressed the computational complexity guarantee of the Learned Index (zeighami2023distribution). This lemma provides a probabilistic guarantee for the error between the output of the PCF and the empirical CDF. Some modifications are required to adapt it to the context of sorting and to attribute it to the probability of bucketing failure, i.e., $\Pr[\exists j,|\bm{c}_{j}|>\delta]$. By appropriately combining the proofs of these two lemmas, Lemma˜3.4 is proved. The exact proof is given in Section˜A.3.
Here, we emphasize that the assumption of this lemma, $0<\sigma_{1}\leq\sigma_{2}<\infty$, is sufficiently reasonable and “mild” as described in (zeighami2023distribution). It asserts that the probability density function $f(x)$ is both bounded and nonzero over its domain $\mathcal{D}$. This class of distributions covers the majority of real-world scenarios because real-world data is commonly derived from bounded and continuous phenomena, e.g., age, grades, and data over a period of time. Empirically, zeighami2023distribution further suggest that the ratio $\sigma_{2}/\sigma_{1}$ tends to remain close to 1 in a wide variety of practical datasets, and even for more challenging cases like OSM, the ratio still appears to remain at most around 20.
Using Lemma˜3.1, Lemma˜3.3, and Lemma˜3.4, we can prove the following theorems.
@@
If $\mathcal{M}_{\mathrm{PCF}}$ is the bucketing method, the worst-case complexity of standard sort is $\mathcal{O}(nU(n))$ (where $U(n)$ is a non-decreasing function), and $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, then the worst-case complexity of PCF Learned Sort is $\mathcal{O}(nU(n)+n\log\log n)$.
@@
When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$ because (i) the PCF is trained in $\mathcal{O}(\alpha+\beta)=\mathcal{O}(n^{3/4})$, and (ii) the total complexity of inference for $n$ elements is $\mathcal{O}(n)$, since the inference is performed in $\mathcal{O}(1)$ per element. Therefore, since $\gamma+1=\mathcal{O}(n)$, the worst-case complexity of standard sort is $\mathcal{O}(nU(n))$, and $\delta=\lfloor n^{3/4}\rfloor$, we can prove the worst-case complexity of PCF Learned Sort is $\mathcal{O}(nU(n)+n\log\log n)$ by Lemma˜3.1. ∎
@@
Let $\sigma_{1}$ and $\sigma_{2}$ be the lower and upper bounds, respectively, of the probability density distribution $f(x)$ in $\mathcal{D}$, and assume that $0<\sigma_{1}\leq\sigma_{2}<\infty$. Then, if $\mathcal{M}_{\mathrm{PCF}}$ is the bucketing method, the expected complexity of standard sort is $\mathcal{O}(n\log n)$, and $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, then the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$.
@@
When $\alpha=\beta=\gamma=\lfloor n^{3/4}\rfloor$, the computational complexity for model-based bucketing is $\mathcal{O}(n)$. Since $K=\Omega(\sqrt{n})$ when $\alpha=\beta=\gamma=\delta=\lfloor n^{3/4}\rfloor$, $K\geq 1$ for sufficiently large $n$, and
$\frac{2n}{\delta}\exp\left\{-\frac{\alpha K}{2\gamma}\left(1-\frac{1}{K}\right)^{2}\right\}=\mathcal{O}(n^{\frac{1}{4}}\exp(-\sqrt{n}))\leq\mathcal{O}\left(\frac{1}{\log n}\right).$ (2)
Given that $\gamma+1=\mathcal{O}(n)$, the expected complexity of standard sort is $\mathcal{O}(n\log n)$, and $\delta=\lfloor n^{3/4}\rfloor$, it follows from Lemma˜3.3 and Lemma˜3.4 that the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$. ∎
Note that the exact value of $\sigma_{1}$ and $\sigma_{2}$ is not required to run PCF Learned Sort since the parameters for this algorithm, i.e., $\alpha$, $\beta$, $\gamma$, and $\delta$, are determined without any prior knowledge. In other words, PCF Learned Sort can sort in expected $\mathcal{O}(n\log\log n)$ complexity as long as $0<\sigma_{1}\leq f(x)\leq\sigma_{2}<\infty$, even if it does not know the exact value of $\sigma_{1}$ and $\sigma_{2}$. If $\sigma_{1}$ and $\sigma_{2}$ do not satisfy the assumption of Theorem˜3.6, the expected complexity of PCF Learned Sort increases to $\mathcal{O}(nU(n)+n\log\log n)$. Such scenarios arise, for instance, in heavy-tailed distributions, datasets with extremely sparse regions ($\sigma_{1}=0$), or cases where the density is highly concentrated at particular values ($\sigma_{2}=\infty$). However, thanks to the worst-case bound in Theorem˜3.5, the complexity never exceeds $\mathcal{O}(nU(n)+n\log\log n)$. Thus, the algorithm remains robust even under distributions that do not fully satisfy the assumption.
Alternative CDF Models.
Our framework is not limited to PCF; it can also accommodate more expressive CDF models. For instance, we can replace PCF with a spline-based model that approximates the CDF by interpolating the empirical distribution at bin boundaries. Concretely, given $\beta$ intervals, we evaluate the empirical CDF at the endpoints of each interval and then construct a piecewise linear spline that connects these values, yielding a continuous and non-decreasing CDF approximation. We refer to the algorithm that applies this spline-based CDF model within our Learned Sort framework as Spline Learned Sort. We show that spline-based models constructed in this way still satisfy the assumptions of Lemma˜3.3, and thus the expected time complexity guarantee of $\mathcal{O}(n\log\log n)$ remains valid. A detailed theoretical proof for the spline-based case is provided in Section˜A.5.
@@ 4 Experiments
In this section, we empirically validate our theoretical results. First, in Section˜4.1, we confirm that the complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ for both synthetic and real datasets. Then, in Section˜4.2, we conduct experiments with various parameter settings and empirically confirm Lemma˜3.4, a lemma that bounds the probability of bucketing failure and plays a crucial role in guaranteeing the expected complexity of PCF Learned Sort. Finally, in Section˜4.3, we present from sorting time measurements.
Setup.
We experimented with synthetic datasets created from the following four distributions: uniform ($\mathsf{min}=0,\mathsf{max}=1$), normal ($\mu=0,\sigma=1$), exponential ($\lambda=1$), lognormal ($\mu=0,\sigma=1$). The input array was generated by independently taking $n$ samples from each distribution. Only the uniform distribution satisfies the theoretical assumptions required for our complexity guarantees. The other distributions violate these assumptions, but we include them to evaluate the empirical robustness of PCF Learned Sort.
We also used the following 16 real datasets, including Chicago [Start, Tot] (chic), NYC [Pickup, Dist, Tot] (nyc), SOF [Humidity, Pressure, Temperature] (sof), Wiki, OSM, Books, Face (sosd-vldb), and Stocks [Volume, Open, Date, Low] (stks). Further dataset details are provided in Appendix B. For each dataset, we randomly sample $n$ elements, shuffle them, and use them as an input array to examine the complexity of the sort algorithms. Since these are real-world datasets, we cannot definitively determine whether they satisfy our theoretical assumptions. However, as shown in the histograms in Figure˜2, Chicago [Start], NYC [Pickup, Tot], SOF [Humidity], and Face appear to distribute values across a relatively dense and continuous domain, aligning well with our assumptions. In contrast, other datasets exhibit long tails or sparse regions, suggesting that our assumptions may not hold. Furthermore, note that for several datasets (Chicago [Start, Tot], NYC [Dist, Tot], SOF [Humidity, Pressure, Temperature], and Stocks [Volume, Open, Date, Low]), the number of unique values is extremely small (less than 3.2% of the total elements, as detailed in Appendix B).
All experiments were run on a Linux machine equipped with an Intel®  Core™  i9-11900H CPU @ 2.50GHz and 62GB of memory. GCC version 9.4.0 was used for compilation, employing the -O3 optimization flag.
4.1 Computational Complexity of PCF Learned Sort